# Learning Program Behavioral Models from Synthesized Input-Output Pairs

### v1

@@ 2.2.2. Masked Tokenization
While our approach works with generic tokenizers, we obtain better results with custom tokenizers adapted to the languages at hand. Our HTMLTokenizer adapted to HTML, for instance, takes an HTML string such as
@@
The result of the tokenization looks like following:
@@
The reconstruction routine is inverse to tokenization, and for many data formats, it can be performed as a simple string concatenation. However, Modelizer provides several implementations for mapped string reconstruction. We support both early and late token instantiations, meaning that abstract tokens will be replaced with real-world values before or after the token concatenation. Furthermore, we support different mapping policies; see Section 3.2 for details of our implementation, including the pseudocode for HTML tokenization.
Full control over the tokenization process allows us to specify the tokenization granularity, impacting the prediction accuracy and model learning time. The vocabulary size grows corresponding to the number of unique tokens in the training set, which impacts the overall number of model parameters. It is useful to combine tokens that sequentially appear in the same context. With this measure, we reduce
@@
How does one implement a domain-specific tokenizer? The simplest way is by to extend the parsing functionality, by deriving a parser from the grammar. In this case, we can reuse the grammar that was used for synthesizing inputs. Still, we cannot infer the output parser this way yet. Once we can automatically and correctly extract output grammars from PUTs, we will be able to infer domain-specific parsers and tokenizers for them.
@@ 2.3. Model Learning
Our Learner is based on a sequence-to-sequence neural machine translation model called Transformer (Vaswani et al., 2017). It learns a program-specific behavior mocking model from selected input-output pairs.
We present the high-level representation of the learning pipeline in LABEL:lst:learn and cover the steps taken in detail in the following paragraphs. Model Learner accepts the input-output data formats as input and loads the corresponding synthesized training sample pairs from the dataset. The loaded samples get automatically split into training, validation, and test sets.
@@
Modelizer automates the hyperparameter tuning process. We have implemented a Hyperparameter Optimizer that tries to automatically find the best combination of configurable parameter values using a small disjoint set of test samples. Such a test set consists of synthesized 10,000 samples. The optimizer is built upon the Optuna (Akiba et al., 2019) framework, which is a hyperparameter optimization framework that can be used for any machine learning model. Parameter search is performed in two phases. The first phase is Model Parameter Search, which looks for the best configuration of the neural network for the fixed number of iterations (Vaswani et al., 2017). Using a fixed learning rate, the optimizer computes the value of the cross entropy loss metric for the different combinations of the Linear Layer size, Embedding size, the number of Attention Heads, and the number of Encoder and Decoder layers. The combination that has the lowest value of cross-entropy criterion is then used in the Learning Rate Search phase, where the starting learning rate (LR) value, as well as weight decay and learning rate adjustment policy, are determined. For our experiments, we have configured the Hyperparameter Optimizer to execute the Model Parameter Search phase for 250 trials and the Learning Rate Search phase for 100 trials.
Then the Learner gets initialized using the discovered model configuration and the model learning process gets executed for the specified number of iterations, called epochs. We use the cross-entropy loss function (Shannon, 1948), which is a generalization of the Logarithmic Loss function for multi-class classification problems, as our training evaluation criterion. It measures the dissimilarity of probability distributions between the expected and predicted labels. Also, we use the Adaptive Moment Estimation with Decoupled Weight Decay Regularization optimization algorithm (AdamW) (Loshchilov and Hutter, 2019) to automatically adjust the model’s parameters to minimize the training loss value. The validation split is used to evaluate the model’s performance on unseen data to prevent overfitting.
With our “lightweight” models, which, for example, require less than 6.3 million parameters to learn the Markdown to HTML conversion behavior of the Pandoc markup converter (MacFarlane, 2006), we achieve high accuracy in mocking program-specific behavior in a short amount of time on a commodity hardware. For the rest of the subjects that we used in the evaluation, the forward models contained 3-to-7.7 million parameters and inverse models required 2.8 to 10.3 million parameters to learn the hypothetical inverse behavior. The training time varies from several minutes to a couple of hours, depending on the subject.
@@ 2.4. Model Deployment
Figure 4. Model Deployment. An example scenario of the Learned Behavior Model predicting Program Input given the Program Output. The model prediction gets validated for correctness in the optional validation phase.
@@
The model Deployment phase, which is summarized in Figure 4, includes the Prediction and optional Validation phases. Depending on the source data type, which can be either program input or output, the respective model, source tokenizer, and target tokenizer get initialized. From this point, model inference operates similarly to the model training phase. The source data gets tokenized (1) and passed to the Program Model (2), which predicts a target token sequence. Then, the predicted sequence gets lifted from token representation to real-world format by the target reconstructor component (3). The translation routine supports an optional max output sequence length parameter. If the framework user does not specify this argument, our default implementation dynamically estimates the maximum output sequence length by applying a 1.25 multiplier to the length of the currently processed input sequence.
Lastly, the predicted data can be passed to an optional Validation phase (4), which requires access to the respective PUT. It operates according to the model’s operation mode. For the behavior-mocking mode (predicting program output from a given program input), predicted output gets compared with computed output, which additionally gets tokenized and is marked as a reference sequence. Using both token sequences, we compute the Levenshtein distance (Levenshtein, 1965) and the BLEU (Papineni et al., 2002) score. Levenshtein distance, also named the edit distance in literature, is a metric for measuring the difference between two sequences, computed as a minimal number of insertions, deletions, and substitutions required to change the predicted sequence into a reference sequence. More on the BLEU score and other sequence similarity metrics will be covered in the next section.
In the Inverse-mocking mode (finding the program input used to compute the given program output), the computed output collected by passing the predicted input to the PUT gets compared with the initial program output. Since the actual program input (ground truth) is unknown, currently we can compute sequence similarity score metrics in cases when the program accepts the predicted input and does not fail while processing it. A failure of course will indicate the incorrectness of the predicted input. It is also challenging to validate the correctness of the prediction in the inverse-mocking mode if the PUT computes multiple valid inputs to the same output.
@@
20 print("BLEU Score:", bleu_score, "Edit Distance:", edit_distance")
@@
22 print("Model predicted", reconstructed")
@@
Such programs have to be modeled by introducing an additional abstraction layer that defines the set of semantically equivalent inputs using formal rules.
We want to demonstrate the results of the model deployment process depicted in LABEL:lst:deploy using the example from Table 1. We used the HTML to Markdown conversion model learned from the Pandoc program, which was trained on 1,000,000 synthesized document pairs. The input HTML string is passed to the model, which predicts the corresponding Markdown string. At first, the input HTML string is tokenized and the content gets replaced with placeholders (Step 1). The model uses the produced input tokens to predict a sequence of Markdown tokens (Step 2). During the reconstruction phase (Step 3) the placeholder tokens in the predicted sequence get replaced with the real-world data. Since the Pandoc program was locally installed on the same machine, which was used to run the HTML to Markdown conversion model, we can validate the correctness of the predicted output by comparing it with the computed output. We measured the BLEU score at 1.0 and the Levenshtein distance at 0.0, which indicates that the predicted output is identical to the computed output.
@@
Input HTML Reconstructed Markdown
@@
Input Tokens Predicted Tokens
@@
The conversion was performed by a model of the Pandoc program trained on 1,000,000 synthesized sequence pairs. BLEU score = 1.0, Levenshtein distance = 0.0
@@ 3. Implementation
Modelizer implements grammar-based input generators using the grammar fuzzers from FuzzingBook(Zeller et al., 2023). While the Fuzzing Book GrammarFuzzer randomly chooses the expansion rule from the set of available rules, its ProbabilisticGrammarFuzzer allows the developer to control the input generation by assigning probabilities to certain expansion rules. In this case, the non-attributed expansion rules will be assigned by equally distributed remaining probabilities.
Such a technique allows us the following:
  • Automatically generating a dataset of syntactically valid inputs.
  • Customizing the feature selection procedure in the sample generation process.
  • Ensuring the representation of specific and corner samples in the data corpus.
@@ 3.1. System Requirements
We implemented the model learning functionality of Modelizer in Python using the PyTorch 2.0 Machine Learning Framework (Paszke et al., 2019). Our framework currently supports the following hardware backends:
  • NVIDIA GPU that supports CUDA 11.7 or higher (Fastest)
  • Apple ARM-based system-on-a-chip with at least 16GB of system memory
  • System with a x64 processor and 32 GB of system memory (Slowest)
While model learning is a highly parallelizable task, and it can benefit from the availability of dedicated hardware accelerators, the rest of the tasks can also be scaled according to the number of available CPU cores. For example, to reduce the total runtime, our generators support the execution of multiple grammar-based fuzzers with the same configuration in parallel. The amount of Fuzzers that got initialized depends on the number of available CPU cores. Modelizer is responsible for distributing the total workload among all fuzzers by assigning every fuzzer with its own budget. Then each fuzzer produces samples in batches until it runs out of the budget. Modelizer systematically collects the generated samples and checks for their uniqueness. If there was no previously unseen sample generated within a single batch or a configurable parameter of failing attempts is reached, the value of the minimum and maximum non-terminals gets automatically incremented. To simplify the usability and applicability of our framework, we include a base class generator that can be extended to support new subjects.
@@ 3.2. Tokenization with Placeholders
Even though our models are trained on synthesized data, they must still be capable of operating with real-world data. However, there can be a difference in the representation between real-world and synthesized data due to the latter optionally including placeholders instead of real-world values, which impacts the implementation of our tokenizers. As a part of our contribution, we supplied the Modelizer with the base implementation and abstract interfaces for the mapped tokenization and mapped reconstruction routines. For every new data type, the framework user can wrap the functionality of unsupervised tokenizers with our mapped tokenization algorithm or implement a format-specific tokenizer in the subclass.
@@ 3.2.1. Abstract Tokenizers
By assuming that for both input and output formats, a suitable parser can be obtained, we implemented AbstractTokenizer—a generic mapped tokenization-reconstruction algorithm that can benefit from format structure knowledge and replications of input content fragments in the output. While it is not a fully functional tokenizer, it provides a high-level generic and extendible implementation of mapped tokenization-reconstruction routines, which assist in mapping real-world content to abstract placeholders and vice-versa. LABEL:lst:tokenize-base contains the fragment of the mapped tokenization routine.
@@
9 def mask_token(self, token: str, pldr: str) -> str:
10 tag = pldr + "_" + str(len(self.mappings[pldr]) + 1)
@@
12 self.mappings[pldr][tag] = token
13 tag = pldr
@@
18 self.mappings[pldr][tag] = token
@@
21 self.mappings[pldr][tag] = token
@@
To be able to correctly map real-world content to abstract placeholders (we use the term mask tokens, which you can see in the pseudocode) whether a parser-tokenizer processes content fragment of data the provided mask_token method must be called. Thus, the framework user must provide the format-specific implementation of the tokenize routine in a subclass. Then algorithm can correctly replace the content with an abstract placeholder during tokenization and successfully recover the original sequence of characters during the reconstruction. During initialization, the tokenizer, can learn the placeholders from the grammar definition automatically or operate with the user-defined collection of placeholders. The proposed algorithm allows reusing the same tokenizer for tokenizing both synthesized and real-world data.
@@ 3.2.2. A HTML Tokenizer
Below we demonstrate a partial implementation of the HTMLTokenizer class, which demonstrates the procedure of masking HTML content by extending the functionality of the AbstractTokenizer class.00footnotetext: The full implementation of HTMLTokenizer, which decomposes data using Python built-in HTMLParser, can be found in our GitHub repository (see Section 7 for details). This repository also provides additional examples of tokenizer implementations that rely on PUT’s functionality to parse inputs or implement a custom parser using regular expressions.
@@
Once the mapped_tokenize method is called, the tokenizer clears all internal buffers and starts processing the given string with the selected Parser. In this implementation, the HTMLParser requires the end-user to implement handler methods to process HTML tags together and document content. Here we present the simplified implementation of the document content handler handle_data. Depending on the operation mode of the tokenizer, which is determined by the mask_mapping flag, the tokenizer either replaces the document content with a matching placeholder or performs a sequence of format-specific string splitting operations, which are abstracted in this example.
@@ 3.2.3. Token Masking Strategies
The token masking algorithm behaves differently according to one of the selected token mapping policy strategies:
  • The SIMPLIFIED policy reduces the vocabulary size of the model by removing identifiers from placeholder tokens.
  • The OPTIMIZING policy, which is the default policy in our implementation, uses a minimal number of placeholder tokens to encode data content.
@@
We use the extended version of the HTML document from Figure 1 to illustrate the effect of using different token masking strategies in tokenization:
@@
When the tokenizer is running in SIMPLIFIED token masking mode, then it removes every unique identifier that had been assigned to a placeholder. It significantly reduces the model’s vocabulary size, but increases the chances of producing errors during the reconstruction phase. The tokenized sequence looks like the following:
@@
The tokenizer tries to use a minimal number of placeholders to encode the real-world data when it is configured to use OPTIMIZING token masking strategy. The repetitive content is encoded by the same placeholder. This is the default operation mode and evaluation experiments were arranged only using this policy. The tokenized sequence looks like the following:

### v2

@@ 2.2.1. Masked Tokenization
While our approach also works with unsupervised tokenizers, such as Google SentencePiece (Kudo and Richardson, 2018), we obtain better results with custom tokenizers adapted for the languages at hand. For instance, our HTMLTokenizer handles an HTML string such as:
@@
The result of the tokenization looks like the following:
@@
Modelizer supports both early and late token instantiations, meaning that abstract tokens will be replaced with real-world values before or after the token concatenation. Furthermore, as well as different mapping policies;. See Section 3.2 for details of our implementation, including the pseudocode for HTML tokenization.
Full control over the tokenization process allows us to control the tokenization granularity, impacting prediction accuracy and model learning time. The vocabulary size grows with to the number of unique tokens in the training set, which impacts the overall number of model parameters. It is worth combining tokens that sequentially appear in the same context. With this measure, we reduce
@@
Modelizer users can implement a domain-specific tokenizer by extending the functionality of parsers that are already available or could be derived from the grammar, for example, with the help of the Earley (Earley, 1970) parser algorithm. In this case, one can reuse the grammar that was used for synthesizing inputs. Once we can automatically and correctly extract output grammars from PUTs, we will be able to infer parsers and domain-specific tokenizers for them. See Section 3.2 for more information on implementing domain-specific tokenizers.
@@ 2.3. Model Learning
Our Learner is based on the Transformer sequence-to-sequence neural machine translation model (Vaswani et al., 2017). It learns a program-specific behavior mocking model from selected input-output pairs. We present the high-level representation of the learning pipeline in LABEL:lst:learn and cover the steps taken in detail in the following paragraphs. The Model Learner accepts the input-output data formats as input and loads the corresponding synthesized training sample pairs from the dataset. The loaded samples get automatically split into training, validation, and test sets.
@@
Modelizer additionally automates the hyperparameter tuning process. We have implemented an Optimizer that tries to automatically find the best combination of configurable parameter values using a small disjoint set of test samples, which, in our case, typically consists of 10,000 synthesized samples. The optimizer is built upon the Optuna (Akiba et al., 2019) framework, which automates hyperparameter search. Parameter search is conducted in two phases: Model Parameter Search and Learning Rate Search. During Model Parameter Search, the best neural network configuration is sought for a fixed number of iterations (Vaswani et al., 2017). While using a fixed learning rate, the Optimizer computes the cross-entropy loss (Shannon, 1948) metric for various combinations of Linear Layer size, Embedding size, number of Attention Heads, and Encoder-Decoder layers. The combination with the lowest cross-entropy value proceeds to the Learning Rate Search phase, where the initial learning rate (LR), weight decay, and learning rate adjustment policy are determined. In our experiments, the Optimizer executes the Model Parameter Search for 250 trials and, afterward, the Learning Rate Search for 100 trials. Automated hyperparameter search facilitates a faster and more effective exploration of the model configuration landscape, which is essential for optimal model performance and training efficiency. It saves the time and effort needed for manual tuning, allowing potential Modelizer users to focus on strategic aspects of program behavior extraction.
After hyperparameter search, the Learner is initialized with the discovered model configuration, and learning runs for the specified epochs. We use the cross-entropy loss function to evaluate training, measuring the difference between expected and predicted labels. The AdamW optimization algorithm (Loshchilov and Hutter, 2019) adjusts model parameters to minimize training loss. The validation split is used to evaluate the model’s performance on unseen data to prevent overfitting. Trained models are “lightweight”, which, for example, require less than 6.3 million parameters to learn the Markdown to HTML conversion behavior of the Pandoc markup converter (MacFarlane, 2006). Meanwhile, we are able to quickly and accurately mock program-specific behavior on commodity hardware. The training time varies from several minutes to a couple of hours, depending on the subject. For the subjects from the evaluation set, the forward models contained 3 to 7.7 million parameters, and inverse models required 2.8 to 10.3 million parameters to learn the hypothetical inverse behavior. Since leveraging the knowledge embedded in pre-trained LLMs could be beneficial for mocking program behavior, we additionally provide an adaptor for interacting with LLMs as an alternative to training task-specific models from scratch. We relied on the Unsloth (Han et al., 2023) framework to support resource-efficient local LLM querying and fine-tuning.
@@ 2.4. Model Deployment
Figure 4. Model Deployment. An example scenario of the Behavior model predicting program input given the program output.
Phases: (1) Input tokenization, (2) Prediction generation, (3) Output reconstruction, (4) Prediction validation.
@@
The Model Deployment phase, summarized in Figure 4, includes the Prediction and optional Validation phases. Depending on the source data type, which can be either program input or output, the respective model, source tokenizer, and target tokenizer get initialized. From this point, model inference operates similarly to the model training phase. The source data gets tokenized (1) and passed to the Program Model (2), which predicts a target token sequence. Then, the target reconstructor component (3) lifts the predicted sequence from token representation to real-world format. The translation routine supports an optional max output sequence length parameter. If the framework user does not specify this argument, our default implementation dynamically estimates the hypothetical maximum output sequence length by applying a 1.25 multiplier to the length of the currently processed input sequence.
Lastly, the predicted data can be passed to an optional Validation phase (4), which requires access to the respective PUT. It aligns with the model’s operation mode. The predicted and computed outputs are tokenized in behavior-mocking mode (predicting output from input), with the latter treated as a reference sequence. Then, we calculate the sequence similarity metrics such as Levenshtein distance (Levenshtein, 1965) and the BLEU (Papineni et al., 2002) score. The Levenshtein distance measures the difference between sequences by counting the number of minimal edits needed to transform the predicted sequence into the reference. The next section will discuss the BLEU score and sequence similarity metrics in more detail.
@@
20 print("BLEU Score:", bleu_score, "Edit Distance:", edit_distance)
@@
22 print("Model predicted", reconstructed)
@@
In the Inverse-mocking mode (finding the program input used to compute the given program output), the computed output collected by passing the predicted input to the PUT gets compared with the initial program output. Since the actual program input (ground truth) is unknown, we can compute sequence similarity score metrics in cases when the program accepts the predicted input and does not fail while processing it. A failure, of course, will indicate the incorrectness of the predicted input. It is also challenging to validate the correctness of the prediction in the inverse-mocking mode if the PUT computes multiple valid inputs to the same output. Such programs have to be modeled by introducing an additional abstraction layer that defines the set of semantically equivalent inputs using formal rules.
We demonstrate the results of model deployment shown in LABEL:lst:deploy using the example from Table 1. In this example, we used the HTML to Markdown conversion model learned from the Pandoc program, which was trained on 1,000,000 synthesized document pairs. First, the input HTML string is tokenized, and the content gets replaced with placeholders (Step 1). The model uses the produced input tokens to predict a sequence of Markdown tokens (Step 2), and placeholders are replaced with real data during output reconstruction (Step 3). Since the Pandoc program was locally installed on the same machine that ran the HTML to Markdown conversion model, we can validate the correctness of the predicted output by comparing it with the computed output. We measured the BLEU score at 1.0 and the Levenshtein distance at 0.0, which indicates that the predicted output is identical to the computed output.
@@
1. Input HTML 4. Reconstructed Markdown
@@
2. Input Tokens 3. Predicted Tokens
@@
The conversion used a Pandoc model trained on 1,000,000 sequence pairs, achieving a BLEU score of 1.0 and a Levenshtein distance of 0.0.
@@ 3. Implementation
Modelizer implements grammar-based input generators using grammar fuzzers from The Fuzzing Book (Zeller et al., 2023). While the Fuzzing Book GrammarFuzzer class randomly chooses the expansion rule from the set of available rules, its ProbabilisticGrammarFuzzer variant allows the developer to control the input generation by assigning probabilities to certain expansion rules. In this case, the non-attributed expansion rules will be assigned by equally distributed remaining probabilities. This allows us to (1) automatically generatea dataset of syntactically valid inputs; (2) customize the feature selection procedure in the sample generation process; and (3) ensure the representation of specific and corner samples in the data corpus.
@@ 3.1. System Requirements
We implemented Modelizer model learning in Python using the PyTorch 2.0 Machine Learning Framework (Paszke et al., 2019). Our framework currently supports the following hardware backends:
  • NVIDIA GPUs with support for CUDA 11.7 or higher (Fastest)
  • Apple Silicon processors with at least 16GB of system memory
  • Systems with a x64 processor and 32 GB of system memory (Slowest)
Model learning is a highly parallelizable task, and it can benefit from the availability of dedicated hardware accelerators; however, the remaining tasks can also be scaled according to the number of available CPU cores. For example, to reduce the total runtime, our generators support the execution of multiple grammar-based fuzzers with the same configuration in parallel. The number of Fuzzers that were initialized depends on the number of available CPU cores. Modelizer is responsible for distributing the total workload among all fuzzers by assigning every fuzzer with its own budget. Then, each fuzzer produces samples in batches until it runs out of budget. Modelizer systematically collects the generated samples and checks for their uniqueness. If there was no previously unseen sample generated within a single batch or a configurable parameter of failing attempts is reached, the value of the minimum and maximum number of non-terminal expansions gets automatically incremented. To simplify the usability and applicability of our framework, we include a base class generator that can be extended to support new subjects.
@@ 3.2. Tokenization with Placeholders
Even though our models are trained on synthesized data, they must still be able to handle real-world data. Since synthesized data may include placeholders instead of actual values, Modelizer provides a base implementation and abstract interfaces for mapped tokenization and reconstruction routines. Users can wrap unsupervised tokenizer functions with our mapped tokenization algorithm or implement a format-specific tokenizer in a subclass for new data types.
@@ 3.2.1. Abstract Tokenizers
Assuming the availability of suitable parsers for input and output formats, we implemented AbstractTokenizer—a generic mapped tokenization-reconstruction algorithm leveraging format structure and content fragment replications in both input and output. While this is not a fully functional tokenizer, it offers a high-level, extendable implementation for mapped tokenization-reconstruction tasks, aiding the mapping of real-world content to abstract placeholders and vice-versa. LABEL:lst:tokenize-base shows a part of the mapped tokenization logic.
To correctly map real content to abstract placeholders (references as mask tokens in the pseudocode), the parser-tokenizer process calls the mask_token method once it handles the data fragment. The developer needs to implement a format-specific tokenize logic in a subclass. Then, Modelizer can replace content fragments with placeholders during tokenization and recover the original sequence during reconstruction. The tokenizer can learn placeholders from the grammar definition or refer to the user-defined ones, allowing it to handle both synthesized and real-world data.
@@
9 def mask_token(self, token: str, placeholder: str) -> str:
10 tag = placeholder + "_" + str(len(self.mappings[placeholder]) + 1)
@@
12 self.mappings[placeholder][tag] = token
13 tag = placeholder
@@
18 self.mappings[placeholder][tag] = token
@@
21 self.mappings[placeholder][tag] = token
@@ 3.2.2. A HTML Tokenizer
HTMLTokenizer class given below demonstrates the procedure of masking HTML content by extending the functionality of the AbstractTokenizer class. Once the mapped_tokenize method is called, the tokenizer clears all internal buffers and starts processing the given string using the selected parser. The HTMLParser module from the Python Standard Library, for instance, requires the user to implement handler methods to process HTML tags and document content (handle_data method). Depending on the operation mode of the tokenizer, which is determined by the mask_mapping flag, the tokenizer either replaces the document content with a matching placeholder or performs a sequence of format-specific string splitting operations, which are abstracted in this example.00footnotetext: The full implementation of HTMLTokenizer, which decomposes data using Python built-in HTMLParser, can be found in our GitHub repository (see Section 7 for details). This repository also provides additional examples of tokenizer implementations that rely on PUT’s functionality to parse inputs or implement a custom parser using regular expressions.
@@ 3.2.3. Token Masking Strategies
The token masking algorithm behaves differently according to the selected token mapping policy strategies:
  • The SIMPLIFIED policy reduces model vocabulary size by removing identifiers from placeholder tokens.
  • The OPTIMIZING policy, our default, uses a minimal number of placeholder tokens to encode data.
@@
We use the extended HTML document from Figure 1 to show the impact of all token masking strategies in tokenization:
@@
When the tokenizer is running in SIMPLIFIED token masking mode, it removes every unique identifier assigned to a placeholder. It significantly reduces the model’s vocabulary size, but increases the chances of producing errors during the reconstruction phase. The tokenized sequence looks like the following:
@@
When configured with the OPTIMIZING token masking strategy, the tokenizer uses a minimal number of placeholders to encode real-world data, reusing the same placeholder for repetitive content. This is the default mode, and it was used for the entire evaluation. The tokenized sequence looks like the following: